# Identification with Orthogonal Basis Functions: Convergence Speed, Asymptotic Bias, and Rate-Optimal Pole Selection

### v1

@@ I Introduction
System identification has long been a focus in the design of control systems, and more broadly in the modeling of dynamical systems [Ljung1998], statistical time series analysis [Madsen2008], and in system realization theory [Picci2015], with early inspirations going back to the works of Zadeh [Zadeh1962] and Kalman [Kalman1965] in the 1960s. Traditional identification methods usually address this problem in the asymptotic sense, where the number of data samples tends to infinity. The central issue is whether the identified mathematical model asymptotically approaches the true physical system, depending on a priori information available about the system and a posteriori data acquired experimentally, and, if so, whether a bias arises between the model and the system. Classical identification algorithms, such as the Ho-Kalman algorithm [ho1966effective] and the multivariable output-error state space (MOESP) approach [Verhaegen1994], are known to be asymptotically unbiased given infinite samples under certain conditions [Ljung1998, Bauer2000, Van1996]. However, these algorithms are typically nonlinear (involving nonlinear operations such as SVD or matrix inversion of the Hankel matrices) with respect to the collected data and may exhibit high sensitivity to noise when only finite-length samples are available [chiuso_ill-conditioning_2004, hachicha2014n4sid], resulting in potential high signal-to-noise ratio requirement and thus high sample complexity. This behavior is also illustrated in our simulation study in Fig. 5.
@@ III Problem Formulation
    $G(z)$ and $H(z)$ are both stricly stable transfer functions; that is, all the poles of $G(z)$ and $H(z)$ are in $\mathbb{D}$. Additionally, $G(z)$ has relative degree at least $1$.
@@
Under Assumptions 1 and 2, the plant transfer function $G(z)$ is uniquely identifiable from closed-loop data with known controller dynamics and injected probing signals. The noise model $H(z)$ is not explicitly identified in this work, and its uniqueness is not relevant to the subsequent analysis.
@@ V Fundamental Lower Bound and Tsuji Pole Selection Algorithm
This section focuses on robustly [^1] One possible approach is to adaptively select $\mu_{k}$ to be as close as possible to the true system poles $\lambda_{j}$ based on the sampled data. However, in system identification, the values of $\lambda_{j}$ are generally unknown. In Section VI, we show that identifying the true system poles becomes exponentially difficult as the system dimension increases. selecting the poles $\mu_{k}$ by minimizing the worst-case identification bias across a class of systems, specifically, those whose poles lie within the pole region [^2] The pole region $\mathcal{D}$ is a complex subset of the open unit disk that contains all true poles [van_den_hof_system_2005, Chapter 11]. $\mathcal{D}$, determined by a priori information. Additionally, we establish a fundamental lower bound on the worst-case identification bias and show that the Tsuji points asymptotically achieve this bound as the number of basis poles $q$ tends to infinity, demonstrating the optimality of the Tsuji pole selection algorithm.
@@ VII Simulations
  • SQP method [bachnas2023advancing]: The min-max pole selection problem (Problem 2) is solved using a multi-start Sequential Quadratic Programming approach implemented via MATLAB’s fmincon, following [bachnas2023advancing, Chapter 3.4.1, Algorithm 1], with pole region $\mathcal{D}=[-0.95,0.95]$.
  • RA method [bachnas2023advancing]: The min-max pole selection problem (Problem 2) is solved using the Randomized Algorithm introduced in [bachnas2023advancing, Chapter 3.4.2, Algorithm 3], based on probabilistic performance verification and bisection search, with pole region $\mathcal{D}=[-0.95,0.95]$.
@@
Figure 2: Asymptotic $\mathcal{H}_{2}$ approximation bias of the nominal system (39) under Gaussian perturbations of the true system poles, visualized using a multi-kernel density estimate (KDE).
@@
The simulation results are shown in Fig. 4. For this high-dimensional system, OBFs with poles selected by the minimax, SQP, RA, and greedy methods exhibit relatively large asymptotic biases, together with increased variability across trials. In contrast, the Tsuji points achieve the smallest asymptotic bias using only $15$ bases and exhibit consistently robust performance, further highlighting the effectiveness of the proposed approach in high-dimensional settings.
Compared with the lower-dimensional example, the advantage of the Tsuji-based pole selection becomes more pronounced in the high-dimensional setting. In low-dimensional systems, small approximation errors may occasionally be observed due to incidental alignment between the selected basis poles and the dominant poles of the true system. As the system dimension increases, such favorable alignment becomes increasingly unlikely, and the consistently small average and worst-case asymptotic biases achieved by the Tsuji points highlight their robustness, especially in high-dimensional systems.
Additionally, we compare the computation time of all pole selection methods as a function of the number of bases $q$, with the results reported in Fig. 6. The initial guess of the Tsuji points exhibits the shortest computation time among all approaches, while the RA method also demonstrates competitive efficiency. The Tsuji points method is slightly slower but remains in the same order of magnitude. Notably, the runtimes of these three methods are all below $10^{-2}$ seconds and do not increase noticeably as $q$ grows. In contrast, the remaining minimax, SQP and greedy methods incur higher computational costs and exhibit a clear increase in runtime for larger $q$. Overall, these results indicate that the proposed Tsuji pole selection achieves a favorable trade-off between approximation accuracy and computational efficiency, particularly in regimes with larger numbers of bases.
@@
Figure 4: Asymptotic $\mathcal{H}_{2}$ approximation bias of the diffusion process under Gaussian perturbations of the true system poles, visualized using a multi-kernel density estimate (KDE).
@@ Appendix B Proof of Theorem 6 and Theorem 8
Equation (4.48) in [van_den_hof_system_2005] provides a pointwise bound on the frequency-response error $|G_{e}(e^{i\omega})|$ on the unit circle. Integrating this pointwise bound over the unit circle immediately yields an $\mathcal{H}_{2}$-type bound that is equivalent to Lemma 15. Related continuous-time variants are also discussed in [Brinsmead2001]. Here, we introduce an alternative proof that utilizes the inner product and the inversion of the Cauchy matrix to derive the discrete-time result.

### v2

@@ I Introduction
System identification has long been a focus in the design of control systems, and more broadly in the modeling of dynamical systems [Ljung1998], statistical time series analysis [Madsen2008], and in system realization theory [Picci2015], with early inspirations going back to the works of Zadeh [Zadeh1962] and Kalman [Kalman1965] in the 1960s. Traditional identification methods usually address this problem in the asymptotic sense, where the number of data samples tends to infinity. The central issue is whether the identified mathematical model asymptotically approaches the true physical system, depending on a priori information available about the system and a posteriori data acquired experimentally, and, if so, whether a bias arises between the model and the system. Classical identification algorithms, such as the Ho-Kalman algorithm [ho1966effective] and the multivariable output-error state space (MOESP) approach [Verhaegen1994], are known to be asymptotically unbiased given infinite samples under certain conditions [Ljung1998, Bauer2000, Van1996]. However, these algorithms are typically nonlinear (involving nonlinear operations such as SVD or matrix inversion of the Hankel matrices) with respect to the collected data and may exhibit high sensitivity to noise when only finite-length samples are available [chiuso_ill-conditioning_2004, hachicha2014n4sid], resulting in potential high signal-to-noise ratio requirement and thus high sample complexity. This behavior is also illustrated in a numerical example in Fig. 6.
@@ III Problem Formulation
    $G(z)$ and $H(z)$ are both strictly stable transfer functions; that is, all the poles of $G(z)$ and $H(z)$ are in $\mathbb{D}$. Additionally, $G(z)$ has relative degree at least $1$.
@@
Under Assumptions 1 and 2, the plant transfer function $G(z)$ is uniquely identifiable from closed-loop data. The noise model $H(z)$ is not explicitly identified in this work, and its uniqueness is not relevant to the subsequent analysis.
@@ V Fundamental Lower Bound and Tsuji Pole Selection Algorithm
This section focuses on robustly [^1] One possible approach is to adaptively select $\mu_{k}$ to be as close as possible to the true system poles $\lambda_{j}$ based on the sampled data. However, in system identification, the values of $\lambda_{j}$ are generally unknown. In Section VI, we show that identifying the true system poles becomes exponentially difficult as the system dimension increases. selecting the poles $\mu_{k}$ by minimizing the worst-case identification bias across a class of systems, specifically, those whose poles lie within the pole region [^2] The pole region $\mathcal{D}$ is a complex subset of the open unit disk that contains all true poles [van_den_hof_system_2005, Chapter 11]. $\mathcal{D}$, determined by a priori information. Additionally, we establish a fundamental lower bound on the worst-case identification bias and show that the Tsuji points asymptotically achieve this bound as the number of OBF poles $q$ tends to infinity, demonstrating the optimality of the Tsuji pole selection algorithm.
@@ VII Simulations
  • SQP method [bachnas2023advancing]: The min-max pole selection problem (Problem 2) is solved using a multi-start Sequential Quadratic Programming (SQP) approach implemented via MATLAB’s fmincon, following [bachnas2023advancing, Chapter 3.4.1, Algorithm 1], with pole region $\mathcal{D}=[-0.95,0.95]$.
  • RA method [bachnas2023advancing]: The min-max pole selection problem (Problem 2) is solved using the Randomized Algorithm (RA) introduced in [bachnas2023advancing, Chapter 3.4.2, Algorithm 3], based on probabilistic performance verification and bisection search, with pole region $\mathcal{D}=[-0.95,0.95]$.
@@
Figure 2: Asymptotic relative $\mathcal{H}_{2}$ approximation bias of the nominal system (39) under Gaussian perturbations of the true system poles, visualized using a multi-kernel density estimate (KDE).
@@
The simulation results are shown in Fig. 4. For this high-dimensional system, OBFs with poles selected by the minimax, SQP, RA, and greedy methods exhibit relatively large asymptotic biases, together with increased variability across trials. In contrast, the Tsuji points achieve the smallest asymptotic bias using only $15$ bases for the $100$-dimensional system and exhibit consistently robust performance, further highlighting the effectiveness of the proposed approach in high-dimensional settings.
Compared with case with the lower-dimensional system, the advantage of the Tsuji-based pole selection becomes more pronounced in the high-dimensional setting. In low-dimensional systems, small approximation errors may occasionally be observed due to incidental alignment between the selected basis poles and the dominant poles of the true system. As the system dimension increases, such favorable alignment becomes increasingly unlikely, and the consistently small average and worst-case asymptotic biases achieved by the Tsuji points highlight their robustness, especially in high-dimensional systems.
Additionally, we compare the computation time of all pole selection methods as a function of the number of bases $q$, with the results reported in Fig. 5. The initial guess of the Tsuji points exhibits the shortest computation time among all approaches, while the RA method also demonstrates competitive efficiency. The Tsuji points method is slightly slower but remains in the same order of magnitude. Notably, the runtimes of these three methods are all below $10^{-2}$ seconds and do not increase noticeably as $q$ grows. In contrast, the remaining minimax, SQP and greedy methods incur higher computational costs and exhibit an increase in runtime for larger $q$. Overall, these results indicate that the proposed Tsuji pole selection achieves a favorable trade-off between approximation accuracy and computational efficiency, particularly in regimes with larger numbers of bases.
@@
Figure 4: Asymptotic relative $\mathcal{H}_{2}$ approximation bias of the diffusion process under Gaussian perturbations of the true system poles, visualized using a multi-kernel density estimate (KDE).
@@ Appendix B Proof of Theorem 6 and Theorem 8
Equation (4.48) in [van_den_hof_system_2005] provides a pointwise bound on the frequency-response error $|G_{e}(e^{i\omega})|$ on the unit circle. Integrating this pointwise bound over the unit circle yields an $\mathcal{H}_{2}$-type bound that is equivalent to Lemma 15. Related continuous-time variants are also discussed in [Brinsmead2001]. Here, we introduce an alternative proof that utilizes the inner product and the inversion of the Cauchy matrix to derive the discrete-time result.